OpenStudy (anonymous):

3 different flavours - cherry, grape and apple.30% buy cherry, 20% buy grapes and 50% buy apple. Cherry buyers buy an average of 1.5 cans. grape buyers at an avereage of 2 and apple buyers at an average of 3.what % of of total sale will cherry be?

OpenStudy (anonymous):

what exactly do u want?

OpenStudy (dumbcow):

total sales = sum of ratio*number of cans --> (.3*1.5) +(.2*2) + (.5*3) = 2.35 so on avg for each customer you expect to sell 2.35 cans of cherry represents the (.3*1.5) \[P = \frac{.3*1.5}{2.35} = 0.1915\] or 19.15%

OpenStudy (anonymous):

Thank u so much!